Cary is 5 years older than Dan. In 6 years the sum of their ages will be 65. Find the age of each man now.
Answer by macston(5194)

D=Dan's age; C=Cary's age=D+5
If the sum of their ages in six years will be 65, the sum is currently 65-12 (subtract 6 years for each)=53 years
D+C=53 Substitute for C
D+D+5=53 Subtract 5 from each side
2D=48 Divide each side by 2
D=24 ANSWER Dan is 24 years old.
C=D+5=24+5=29 ANSWER 2: Cary is 29 years old
CHECK
(D+6)+(C+6)=65
24+6+29+6=65
65=65